OpenStudy (idealist10):

Find all solutions of y'-xy=(x^3)(y^3).

OpenStudy (dls):

Its Bernoulli. Get it into the form of this. \[\LARGE \frac{dy}{dx} + P(y) = Q\]

OpenStudy (dls):

Make sure that Q should be solely a function of x.

OpenStudy (idealist10):

So do I divided the whole equation by y^3?

OpenStudy (dls):

yep!

OpenStudy (idealist10):

Then what do I do?

OpenStudy (rational):

sibstitute v = 1/y^2

OpenStudy (dls):

\[\Large \frac{y'}{y^3}-\frac{x}{y^2}=(x^3)\] Let v=1/y^2

OpenStudy (dls):

rationalllllllllllllllll

OpenStudy (idealist10):

Let me work it out and I'll tell you what I got.

OpenStudy (dls):

sure!

OpenStudy (idealist10):

Since v=1/y^2, so y=1/sqrt(v), right?